OpenStudy (anonymous):

i need help on a question. Jessica deposits $300 into a savings account that pays a simple interest rate of 6%. How much money will Jessica have in her account at the end of six months?

OpenStudy (anonymous):

can anyone help me

OpenStudy (anonymous):

6% annually?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

if 6% is the annual rate, then the bank gives 6$ interest for 100$ deposit in 12 months. so for 6 months, the bank will give 6/12 * 6 $ interest= 3$ interest. so the interest is 3% for 6 months. so how much will be the total money after 6 months? 300(1+3/100) $=309 $

OpenStudy (anonymous):

ahh thank you can u help me with how you found tht answer??

OpenStudy (anonymous):

i already showed you how i got it.. is there anything you did not understand?

OpenStudy (shamim):

or we can use the formula I=Pnr

OpenStudy (shamim):

I=interest money

OpenStudy (shamim):

P=deposit

OpenStudy (shamim):

n=6 months=0.5 year

OpenStudy (shamim):

r= interest rate=3%=3/100=0.03